OpenStudy (anonymous):

(-9,-6) perpendicular to x=3

OpenStudy (anonymous):

x=3 is a vertical line, which has no slope a perpendicular line will have a slope of 0 because it will be a horizontal line y=mx+b fill in what you know -6=-9(0)+b -6=b y=ox-6 or y=-6

OpenStudy (anonymous):

i have another question F(x)=5x^2-3x+6 Find f(k-1)